National Accounts and Price Indices Consider an economy that only produces three types of goods and services: bread, tractors, and haircuts. The following table shows how many of these goods were produced in three consecutive years and at what price they were sold year bread tractors haircuts units sold price units sold price units sold price 2008 2009 2010 4,950,000 5,000,000 5,050,000 $1.00 $1.00 $1.05 95 100 95 $45,000 $50,000 $45,000 96,000 100,000 104,000 $50 $50 $55

According to the statistics agency of this country, a typical household consumes 90 units of bread and 2 haircuts per vear. f) Calculate the CPI implied by this goods bundle for the three years. What are the CPI inflation rates? g)Which of the growth rates you have calculated (or what combination thereof) best describes... i. the change of economic activity, the loss of value of the currency as felt by typical consumers, afford to buy, and economy? iii. the change of the number of consumption goods a typical household can iv. the change of the total dollar value of all transactions that take place in the Briefly explain your answers (max. one sentence each)
Answer #1

f) If we assume base year = 2008

Consumer price index for year 2008 = 100

Consumption of bread by a typical household = 90 units and consumption of haircuts by a typical household = 2 and therefore, we can show weights and price schedule of two goods bread and haircuts in the following table:

Year

Bread

Weight= Single household consumption/total household consumption

Price

Haircuts

Weight

Price

2008

0.0000181

$1

0.000020

$50

2009

0.000018

$1

0.00002

$50

2009

0.0000178

$1.05

0.0000192

$55

The Consumer price index for year 2009 = Weighted average prices of bread and hair cuts in 2009/weighted average prices of bread and hair cuts in 2008*100= (0.0000181*1+0.00002*50)/(0.000018*1+0.00002*50)*100 =100

The consumer price index for year 2010 = Weighted average price of bread and hair cuts in 2010/weighted average prices of bread and hair cuts in 2008 = (0.0000178*1.05+0.0000192*55)/(0.000018*1+0.00002*50)*100 = 105.568

CPI inflation rate for year 2009 = (CPI for 2009-CPI for 2008)/CPI for 2008*100= 0

CPI inflation rate for year 2010= (CPI for 2010-CPI for 2008)/CPI for 2008*100=(105.568-100)/100*100= 5.568%

g) CPI inflation rate for year 2009 best describes change of the total dollar value of all transactions that take place in an economy.

i) The CPI growth or inflation rate of 2009 best describes change in economic activity as it reflects actual increase in total consumption of breads and haircuts of households.

ii) The CPI inflation of 2010 reflects loss of currency value for typical consumers.

iii) The CPI inflation of year 2010 reflects change of number of consumption of goods by a typical household in view of rise in prices.

iv) The CPI inflation of year 2009 reflects change of the total dollar value of all transactions that take place in an economy.